# Vanishing of Conserved Charges in Cotton Gravity

### v1

@@ II.1 Constructing the conserved current
To construct the conserved charges of the Cotton gravity, let us recall some of the properties of the Cotton tensor [18].
@@
Already at this stage, an astute reader will recognize that the conserved current structure of this theory signals trouble in the sense that the current is obtained from the divergence of the Cotton tensor, i.e. the tensor used in defining the field equations. So in a vacuum, far outside the sources, the current is identically zero and this will yield zero conserved charges for all solutions. Recall that in Einstein’s theory, the situation is quite different: one has $J^{\mu}=\xi_{\nu}G^{\nu\mu}=\nabla_{\nu}{\mathcal{F}^{\nu\mu}(\xi,h)}$ where $\mathcal{F}^{\nu\mu}$ is antisymmetric and given explicitly in terms of the deviation of the metric ($h$) from the vacuum and the Killing vector. See equations (28) and (29) in [16].
@@
Following the steps given in Section III of [19] verbatim, one can build the conserved charges for each Killing vector field. To do so, let $\bar{g}_{\mu\nu}$ be the background metric on the manifold which is considered to be the vacuum with zero conserved charges as noted above. We assume that there is at least one Killing vector $\bar{\xi}^{\mu}$ that the vacuum possesses, and let $\bar{\mathcal{M}}$ denote this geometry and $h_{\mu\nu}$ denote a symmetric field living on it in such a way that it decays sufficiently fast to make the relevant integrals finite. Then integrating $\partial_{\mu}\mathcal{J}^{\mu}=0$ over the manifold, and using Stokes’ theorem, one has
@@
which means we can use any $\bar{\Sigma}_{t}$ at time $t$, hence the name conserved charge. In (20), we can apply the Stokes’ theorem once again to get
@@ III CONCLUSIONS
We have shown that the Cotton gravity theory has identically zero conserved charges for all of its solutions as long as there is no matter at the spatial boundary of spacetime. This shows that the theory is unphysical in the sense that the vacuum without a black hole is degenerate with a spacetime with many black holes. Hence the theory does not have a viable ground state. For example, the parameter $m$ in the spherically symmetric solution given in [2] cannot be interpreted as the conserved mass. Let us also note that there could be another, albeit unlikely, explanation for this vanishing conserved-charge result: in pure Weyl-square theory in the metric formulation, a zero energy theorem for all solutions was interpreted as confinement of energy in [17]. This was later reanalyzed and disputed within the Killing charge construction in [15] where a section was devoted to possible zero energy models. Weyl gravity and Cotton gravity are related in the sense that the former involves the usual metric variation of the Weyl-square action, while the latter involves a variation with respect to the Christoffel connection [2].
@@ Acknowledgements.
We dedicate this work to the memory of Prof. Dr. I. Ferit Oktem (1928-2024), one of the first researchers to write papers on relativity from Turkey. We thank Metin Gurses for useful discussions. E. Altas is supported by the TUBITAK Grant No. 123F353.

### v2

@@ II.1 Constructing the conserved current
To construct the conserved charges of Cotton gravity, let us recall some of the properties of the Cotton tensor [18].
@@
Already at this stage, an astute reader will recognize that the conserved current structure of this theory signals trouble in the sense that the current is obtained from the divergence of the Cotton tensor, i.e. the tensor used in defining the field equations. So in a vacuum, far outside the sources, the current is identically zero and this will yield zero conserved charges for all solutions. Recall that in Einstein’s theory the situation is quite different: one has $J^{\mu}=\xi_{\nu}G^{\nu\mu}=\nabla_{\nu}{\mathcal{F}^{\nu\mu}(\xi,h)}$ where $\mathcal{F}^{\nu\mu}$ is antisymmetric and given explicitly in terms of the deviation of the metric ($h$) from the vacuum and the Killing vector. See equations (28) and (29) in [16].
@@
Following the steps given in Section III of [19] verbatim, one can build the conserved charges for each Killing vector field. To do so, let $\bar{g}_{\mu\nu}$ be the background metric on the manifold that is considered to be the vacuum with zero conserved charges as noted above. We assume that there is at least one Killing vector $\bar{\xi}^{\mu}$ that the vacuum possesses, and let $\bar{\mathcal{M}}$ denote this geometry and $h_{\mu\nu}$ denote a symmetric field living on it in such a way that it decays sufficiently fast to make the relevant integrals finite. Then integrating $\partial_{\mu}\mathcal{J}^{\mu}=0$ over the manifold, and using Stokes’ theorem, one has
@@
which means we can use any $\bar{\Sigma}_{t}$ at time $t$, hence the name conserved charge. In (20), we can apply Stokes’ theorem once again to get
@@ III CONCLUSIONS
We have shown that Cotton gravity theory has identically zero conserved charges for all of its solutions as long as there is no matter at the spatial boundary of spacetime. This shows that the theory is unphysical in the sense that a vacuum without a black hole is degenerate with a space-time with many black holes. Hence, the theory does not have a viable ground state. For example, the parameter $m$ in the spherically symmetric solution given in [2] cannot be interpreted as the conserved mass. Let us also note that there could be another, albeit unlikely, explanation for this vanishing conserved-charge result: In pure Weyl square theory in the metric formulation, a zero energy theorem for all solutions was interpreted as the confinement of energy in [17]. This was later reanalyzed and disputed within the Killing charge construction in [15] where a section was devoted to possible zero-energy models. Weyl gravity and Cotton gravity are related in the sense that the former involves the usual metric variation of the Weyl square action, while the latter involves a variation with respect to the Christoffel connection [2].
@@ Acknowledgements.
We dedicate this work to the memory of Prof. Dr. I. Ferit Oktem (1928-2024), one of the first researchers to write papers on relativity from Turkey. The authors thank Metin Gurses for useful discussions. E. Altas is supported by TUBITAK Grant No. 123F353.